Line passing through (5,-6) and (-9,4) is perpendicular to a line with a slope 14/5. Find the value of y
Answer by checkley71(8403)

A LINE PERPENDICULAR TO A LINE WITH SLOPE=14/5 HAS A SLOPE=-5/14.
-6=-5/14*5+b
-6=-25/14+b
b=-6+25/14
b=(-6*14+25)/14
b=-59/14 ANSWER FOR THE Y INTERCEPT.
THUS THE EQUATION FIOR THIS LINE IS:
Y=-5X/14-59/14